# Turbulence-induced anti-Stokes flow: experiments and theory

### v1

@@ 4 Theoretical considerations
Throughout the theory sections, we assume that there is a negligible mean pressure gradient in the flow (apart from the local pressure gradient associated with a wave group) and that the mean flow after averaging over wave and turbulent motion (both time and ensemble average) lies in the streamwise-vertical plane, i.e. $\bar{v}=0$ and all averaged quantities are presumed independent of the spanwise coordinate $y$. At the time and length scales we consider, the Coriolis force is irrelevant.
@@
${\boldsymbol{u}}(\mathbf{x},t)=\bar{{\boldsymbol{u}}}(\mathbf{x})+\tilde{{\boldsymbol{u}}}(\mathbf{x},t)+{\boldsymbol{u}}^{\prime}(\mathbf{x},t)\equiv\check{{\boldsymbol{u}}}(\mathbf{x})+\tilde{{\boldsymbol{u}}}(\mathbf{x},t),$ (7)
where $\tilde{{\boldsymbol{u}}}$ and ${\boldsymbol{u}}^{\prime}$ are the contributions from waves and turbulence, respectively, and the Eulerian-mean flow $\bar{{\boldsymbol{u}}}$ changes slowly compared to the period and wavelength of an individual wave. The wave-filtered (Eulerian, turbulent) current is denoted $\check{{\boldsymbol{u}}}$. Performing the decomposition in practice is non-trivial when the waves are not monochromatic and/or there is no clear scale separation in time or space. We tried and evaluated two different methods for separations — the widely used ‘phase-conditioned averaging’ (PhCA) and proper orthogonal decomposition (POD) — and found the latter to perform better. A detailed comparison is found in Appendix C.
@@ 4.1 Irrotational Eulerian return flow
The return flow is found beneath a passing wave group, following the group and tending rapidly to zero ahead of and behind the group. For this reason, the velocity measurements in cases 1.A1.D  which are taken before and after the passage of the wave group, are expected to see negligible influence from such flow. Cases 2.A3.B, on the other hand, occur under regular waves — in practice a very long wave group. Since the length of the “group” is far longer than the depth of the channel, the current will be approximately depth-uniform and, in the system following the mean current without waves, satisfying Lagrangian mass conservation, i.e.,
$u_{\mathrm{rf}}=-\frac{1}{h}\int_{-h}^{0}u_{s}(z;h)\mathrm{d}z\approx-\frac{1}{h}\int_{-\infty}^{0}u_{s}(z)\mathrm{d}z=-\frac{\epsilon_{0}^{2}c(k_{0})}{2k_{0}h},$ (8)
@@ 4.2 Statistical theory of Eulerian flow generation
Pearson (2018) argued that when waves appear in a turbulent field, the combined flow will undergo a transition to a new statistically steady state. We will tailor his argument to our current application and discuss consequences in the context of our experimental observations. Momentum conservation, expressed through the time-averaged Navier–Stokes equations or the Craik–Leibovich equations as appropriate, implies that an Eulerian-mean current will manifest during the ‘spin-up’ period, before a quasi-steady state is reached (in our case slowly decaying due to dissipation).
@@ 4.2.1 Turbulent statistics in the transition period
While quiescent-water irrotational waves and turbulence may each be steady in isolation (except for their decay due to dissipation), together they form a system in disequilibrium which will go through a transient change to a new quasi-steady state. It is similar to the model of ‘spin-up from rest’ employed in theory for Langmuir turbulence (e.g., McWilliams et al., 1997). Revisiting and, to some extent, adapting the arguments of Pearson (2018) sheds light both on the early-stage ‘spin-up’ stage and the final situation.
We first assume that a triple decomposition of the turbulent and wavy flow can be performed according to (7) (by no means a trivial point, an intricacy we shall return to later) and that all averaged quantities vary slowly as a function of $x$ over a wavelength so that their derivative can be neglected to leading order. Moreover, assume that any changes in the mean flow develop slowly compared to a wave period, which is reasonable since we see in figure 3 that the final stage appears to be far from reached in cases 1.A1.D after the waves and turbulence have interacted for, effectively, $T_{\mathrm{int}}\sim 7$ wave periods. The wave field, and consequently the Stokes drift, changes slowly, so $\partial_{t}{\boldsymbol{u}}_{s}$ is negligible. The flow is assumed to be unforced, without, e.g., wind stress, and there is no influence from buoyancy or the Coriolis force. The details of the dissipation are not important to the argument, so we shall not consider them beyond assuming that viscous decay is slow compared to the mean-flow effects of wave-turbulence interactions (which is reasonable in light of the observations by Jooss et al., 2021), so that a quasi-steady state can be reached. The final assumption is that $\bar{{\boldsymbol{u}}}$ and ${\boldsymbol{u}}_{s}$ are both oriented along the $x$-axis and vary slowly except with coordinate $z$.
@@
$\partial_{t}\check{{\boldsymbol{u}}}+(\check{{\boldsymbol{u}}}\cdot\nabla)\check{{\boldsymbol{u}}}={\boldsymbol{u}}_{s}\times\check{\boldsymbol{\omega}}-\nabla(\check{p}+{\textstyle\frac{1}{2}}u_{s}^{2}+{\boldsymbol{u}}_{s}\cdot\check{{\boldsymbol{u}}})+\text{diffusion,}$, (9)
where $\check{\boldsymbol{\omega}}=\nabla\times\check{{\boldsymbol{u}}}$ denotes the (Eulerian) vorticity field. Taking the $x$-component and ignoring diffusion gives
@@
Equation (11) has two important consequences. First, that an inhomogeneous field of turbulence will drive a mean current for as long as a vertical gradient of the shear stress exists (and dominates over viscous forces). Second, when a steady state has been reached, i.e., $\partial_{t}\bar{u}=0$, then $\overline{u^{\prime}w^{\prime}}$ is approximately constant with respect to depth (see also Pearson, 2018, for more discussion), or, more precisely, its vertical gradient is balanced by viscous diffusion. Physically, $\overline{u^{\prime}w^{\prime}}\neq 0$ represents a vertical redistribution of streamwise momentum, allowing ${\boldsymbol{u}}={\boldsymbol{u}}(z)$ to transition from its initial value to a presumed final steady-state value, so when the transition period is ended, this shear stress should thus be small compared to TKE. Equation (11) paves the way for further analysis of the ‘spin-up’ of the Eulerian-mean current, because the right-hand side can be related to the underlying physical process using a model based on RDT.
@@ 4.3 Statistics in the quasi-equilibrium state
where we have employed the chain rule and the fact that $\nabla\cdot{\boldsymbol{u}}^{\prime}=0$. The diffusion term contains viscous dissipation and turbulent pressure fluctuations, both of which may reasonably be assumed to be small in an oceanographic setting (Harcourt, 2015; Pearson, 2018) (unfortunately we cannot directly ascertain how accurate this assumption is; see also Pearson et al., 2019). We should bear in mind that this assumption is questionable for our experiment, where waves are shorter and turbulence levels considerably higher than in typical field settings. This is especially true for cases 2.A2 and 2.B2. For instance, angular diffusion of the waves can be highly significant when the waves are not long compared to the turbulent integral scale as found by Smeltzer et al. (2023) which would cause a spatially variable Stokes drift. The approximation is more justified the larger the Stokes drift ($\sim\epsilon_{0}^{2}c(k_{0})$) is compared to turbulent velocities ($\sim u_{\infty}$). In cases 2.A2 and 2.B2 the turbulence is strong and the waves short, and turbulent diffusion and pressure correlation terms may not be negligible compared to the Stokes drift contributions even at surface level. Unfortunately, we are not in a position to quantify the diffusive terms in (12) from our data.
@@
which is to say that an Eulerian-mean current must have been created with the opposite sign compared to Stokes drift.
@@
with ${\boldsymbol{u}}_{s}$ and $\bar{{\boldsymbol{u}}}_{h}$ lying in the horizontal plane. We emphasise that if a constant, depth-uniform current is initially present (as in our experiment), the current $\bar{{\boldsymbol{u}}}$ in (13) and (14) is the change in current due to wave-current interaction, or alternatively the current measured in the reference system following the original Eulerian-mean flow.
Beneath a free surface, turbulence is not isotropic, yet $\overline{u^{\prime}u^{\prime}}$ and $\overline{w^{\prime}w^{\prime}}$ can be expected to be of the same order of magnitude except very near the surface where the vertical velocity tends to zero. It is worth noticing that (13) is not a perfect cancellation between Eulerian flow and Stokes drift as suggested by the experimental (re)analysis of Monismith et al. (2007), but when turbulence is close to isotropic, the remaining Lagrangian-mean current could be difficult to distinguish from zero in an experiment. On the basis of available evidence it seems a reasonable conjecture that the wave-turbulence-generated anti-Stokes flow explains, at least partly, the surprising cancellation.
@@ 4.3.1 Comparison with experiment
Figure 6: Test of equation (13) for cases 2.A to 3.B. The theoretical Stokes drift gradient $\mathrm{d}u_{s}/\mathrm{d}z$ (the derivative of equation 1) is shown as dashed lines of corresponding colour. Separation of turbulence from waves was performed with proper orthogonal decomposition, discussed in Appendix C.
In order to test (13), it is necessary to extract the variances $\overline{u^{\prime}u^{\prime}}$ and $\overline{w^{\prime}w^{\prime}}$ from the wave-turbulence PIV data. Accomplishing this with the best possible accuracy is paramount in order to estimate turbulent second-order moments, because the wave velocities far exceed those of the turbulence near the surface, so even a small percentage of the wave motion erroneously identified as turbulence could lead to considerable overestimates of $\overline{u^{\prime}u^{\prime}}$ and $\overline{u^{\prime}w^{\prime}}$ (less so $\overline{u^{\prime}w^{\prime}}$ due to the phase difference between $\tilde{u}$ and $\tilde{w}$).
@@
The two sides of (13) are compared in figure 6 for cases 2.A to 3.B, the cases whose measurements were taken in the streamwise-vertical plane so that waves and turbulence can be separated. The POD procedure was employed for triple decomposition to evaluate turbulent second moments. The dashed lines show the vertical derivative of the Stokes drift, $2\sqrt{gk_{0}}\epsilon_{0}^{2}\exp(2k_{0}z)$, while the solid line is the measured value of $-(\overline{w^{\prime}w^{\prime}}/\overline{u^{\prime}u^{\prime}})\mathrm{d}\bar{u}/\mathrm{d}z$, which is hypothesised to equal the Stokes drift gradient in the quasi-steady state. For completeness, the Reynolds stresses themselves are plotted in Appendix D.
@@
Given the highly irregular and unsteady nature of our strongly turbulent flow, the agreement is striking in many of the cases. As argued above, the assumptions behind (13) are expected to be reasonable when Stokes drift is sufficiently high compared to turbulent velocities, i.e. $k_{0}z\gtrsim 2$ and sufficiently large $\epsilon_{0}$, and subject to corrections if diffusion levels are high. The relatively poor fit for cases 2.A2 and 2.B2 is therefore not very surprising for the latter reason. The very low values of $\epsilon_{0}$ in cases 2.A1.1 and 2.A2.1 is a likely reason why (13) is unlikely to be reasonable for these cases — again the Stokes drift term in (25) may not dominate over terms which are neglected. Indeed, the good performance in case 3.A.1 is more striking, but tallies with the fact that TKE is only half that of case 2.B and less than a quarter of that of case 2.A, with correspondingly less diffusion.
We hesitate to offer an explanation for the surprising fact that the experimental curves in panel figure 6 appear to curve back near the surface, corresponding to a similar behaviour of the fraction $\overline{u^{\prime}u^{\prime}}/\overline{w^{\prime}w^{\prime}}$ (see also figure 12(i)), yet we remind the reader of the difficulty of separating waves and turbulence near the surface for these cases; we cannot rule out that the effect is spurious.
@@ 4.4 A model based on RDT
The early onset of the (change in the) Eulerian current is governed approximately by (11), and we will use RDT to study how its right-hand side—that is, the Reynolds stress $\overline{u^{\prime}w^{\prime}}$—depends on depth $z$ and time while changes are still moderate. The theoretical predictions will be compared to the observations made for cases 1.A1.D, shown in figure 3(b). Teixeira & Belcher (2002) showed using RDT that a shear stress that varies in the vertical is generated by the passage of a progressive monochromatic surface wave over isotropic turbulence. We will show next that something similar occurs for a finite wave group, such as considered in the experiments of the present study. Since the effect we consider is inviscid in nature, we will neglect viscosity in the following.
@@
RDT is a theory first proposed by Batchelor & Proudman (1954) where the straining of turbulence due to the distortions from the surrounding flow is assumed to dominate over that due to turbulence acting on itself, so that the term $({\boldsymbol{u}}^{\prime}\cdot\nabla){\boldsymbol{u}}^{\prime}$ is negligible in the Navier–Stokes equation, yielding a linearised theory. The RDT approach is formally valid whenever the distortions applied to turbulence are sudden. This amounts to assuming that the distortions (for example, mean-flow gradients) are applied over a time shorter than an eddy turn-over time. The spectral formulation of RDT also requires that there is a spatial scale separation between the turbulence and the mean flow. In the present wave-associated mean flow, this corresponds to $\lambda\gg L$ where $\lambda$ is the wavelength of the waves and $L$ is the integral length scale of the turbulence (cf. Teixeira & Belcher (2002)). Both of these criteria are reasonably well satisfied in the experiments 1.A1.D where the relevant turbulent lengthscale, $L_{x}^{x}$ is at most half a wavelength. Visual inspection of the velocity field clearly shows that the typical coherent eddies are much smaller than $\lambda_{0}$. The values of $L_{x}^{x}$ and details of its calculation with the method of Mora & Obligado (2020) were reported in Smeltzer et al. (2023), and are quoted in Table 3. In practice, RDT is known to provide useful results even when these conditions are not strictly fulfilled (e.g. Hunt & Carruthers, 1990; Mann, 1994; Cambon & Scott, 1999).
@@
In the present case, the effect that is essential in order to explain the generation of a Eulerian-mean current is not related to the individual wave oscillations but rather to the systematic tilting and stretching of the vorticity of the turbulence by the Stokes drift of the wave. Therefore, we adopt a linearised version of the Craik–Leibovich equation (9) whose key term is the ‘vortex force’ ${\boldsymbol{u}}_{s}\times\check{\boldsymbol{\omega}}$. It should be noted that these equations represent the effect of the Stokes drift of an irrotational surface wave; the rotational correction to the wave motion due to the small change in Eulerian current (see Ellingsen, 2016) is neglected. The corresponding vorticity equation, which will be used in RDT, is
@@
In (15), ${\boldsymbol{\omega}}^{\prime}=\nabla\times{\boldsymbol{u}}^{\prime}$ is the turbulent vorticity. As before, ${\boldsymbol{u}}_{s}$ is orientated along the $x$-axis and is independent of $x,y$ and $t$. The mean background current is assumed to be initially depth uniform, hence it has only trivial effect on the flow system, and we set it to zero by a change of coordinate system.
Let us first consider the turbulence away from the air-water interface, which to a first approximation can be considered homogeneous and isotropic, being denoted by the superscript $(H)$. In this section, it is useful to let the subscript $i=1,2,3$ denote the component of a vector or tensor along directions $x,y$ and $z$, respectively. The turbulent, homogeneous (or bulk) vorticity ${\boldsymbol{\omega}}^{\prime(H)}$ is expressed as a 3D Fourier integral, as
@@
where the hat denotes a Fourier transformed turbulent perturbation quantity and ${\boldsymbol{\kappa}}=(\kappa_{1},\kappa_{2},\kappa_{3})$ is the wavenumber vector. In RDT, the leading cause of change for the turbulence is presumed to be the Lagrangian motion of turbulence-containing parcels of fluid due to the larger-scale surrounding motion.
@@
$\frac{\mathrm{d}\hat{\omega}^{(H)}_{i}}{\mathrm{d}t}$ $=\frac{\partial({\boldsymbol{u}}_{s})_{i}}{\partial x_{j}}\hat{\omega}^{(H)}_{j},$ (17)
@@
Since $({\boldsymbol{u}}_{s})_{i}=u_{s}\delta_{1i}$, (17)-(18) reduce to
@@
where the subscript ‘00’ applied to a variable denotes its value at the initial time, before the turbulence is distorted. It is convenient to define
@@
where $\varepsilon_{ijl}$ is the Levi–Civita permutation symbol and $\kappa=|{\boldsymbol{\kappa}}|$.
We wish to related the turbulent velocity after distortion by the Stokes drift to the velocity of the initial undistorted turbulence. The Fourier transform of the velocity of the homogeneous turbulence after distortion by the Stokes drift can be related to the corresponding Fourier transform of the vorticity using (23). In turn, the vorticity of the homogeneous turbulence after distortion can be related to the initial turbulent vorticity using (20a)-(20b). Finally, the initial turbulent vorticity can be related to the initial turbulent velocity through

### v2

@@ 4 Theoretical considerations
Throughout the theory sections, we assume that there is a negligible mean pressure gradient in the flow (apart from the local pressure gradient associated with a wave group) and that the mean flow after averaging over wave and turbulent motion (both time and ensemble average) lies in the streamwise–vertical plane, i.e. $\bar{{\boldsymbol{u}}}=(U_{0},0,0)$, i.e., $\bar{v}=\bar{w}=0$, and all averaged quantities are presumed independent of the spanwise coordinate $y$. At the time and length scales we consider, the Coriolis force is irrelevant.
@@
${\boldsymbol{u}}(\mathbf{x},t)=\bar{{\boldsymbol{u}}}(\mathbf{x})+\tilde{{\boldsymbol{u}}}(\mathbf{x},t)+{\boldsymbol{u}}^{\prime}(\mathbf{x},t)\equiv\check{{\boldsymbol{u}}}(\mathbf{x},t)+\tilde{{\boldsymbol{u}}}(\mathbf{x},t),$ (7)
where $\tilde{{\boldsymbol{u}}}$ and $\boldsymbol{u}$$^{\prime}$ are the contributions from waves and turbulence, respectively, and the Eulerian-mean flow $\bar{{\boldsymbol{u}}}$ changes slowly compared with the period and wavelength of an individual wave. The wave-filtered (Eulerian, turbulent) current is denoted $\check{{\boldsymbol{u}}}=\bar{{\boldsymbol{u}}}+{\boldsymbol{u}}^{\prime}$. Performing the decomposition in practice is non-trivial when the waves are not monochromatic and/or there is no clear scale separation in time or space. We tried and evaluated two different methods for separations—the widely used ‘phase-conditioned averaging’ (PhCA) and proper orthogonal decomposition (POD)—and found the latter to perform better. A detailed comparison is found in Appendix C.
@@ 4.1 Irrotational Eulerian return flow
The return flow is found beneath a passing wave group, following the group and tending rapidly to zero ahead of and behind the group. For this reason, velocity measurements in cases 1.A1.D  which are taken before and after the passage of the wave group, are expected to see negligible influence from such flow. Cases 2.A3.B, on the other hand, occur under regular waves – in practice a very long wave group. Since the length of the “group” is far longer than the depth of the channel, the current will be approximately depth-uniform and, in the system following the mean current without waves, satisfying Lagrangian mass conservation, i.e.,
$u_{\mathrm{rf}}=-\frac{1}{h}\int_{-h}^{0}u_{s}(z;h)\mathrm{d}z\approx-\frac{1}{h}\int_{-\infty}^{0}u_{s}(z)\mathrm{d}z=-\frac{k_{0}a^{2}}{2h}c(k_{0}),$ (8)
@@ 4.2 Statistical theory of Eulerian flow generation
Pearson (2018) argued that, when waves appear in a turbulent field, the combined flow will undergo a transition to a new statistically steady state. We will tailor his argument to our current application and discuss consequences in the context of our experimental observations. Momentum conservation, expressed through the time-averaged Navier–Stokes equations or the Craik–Leibovich equations as appropriate, implies that an Eulerian-mean current will manifest during the ‘spin-up’ period, before a quasi-steady state is reached (in our case slowly decaying due to dissipation).
@@ 4.2.1 Turbulent statistics in the transition period
While quiescent-water irrotational waves and turbulence may each be steady in isolation (except for their decay due to dissipation), together they form a system in disequilibrium which will go through a transient change to a new quasi-steady state. This is similar to the model of ‘spin-up from rest’ employed in theory for Langmuir turbulence (e.g., McWilliams et al., 1997). Revisiting and, to some extent, adapting the arguments of Pearson (2018) sheds light both on the early-stage ‘spin-up’ stage and the final situation. We first assume that a triple decomposition of the turbulent and wavy flow can be performed according to (7) (by no means a trivial point, an intricacy we shall return to later) and that all averaged quantities vary slowly as a function of $x$ over a wavelength so that their derivative can be neglected to leading order. Moreover, we assume that any changes in the mean flow develop slowly compared with a wave period. This is reasonable because we have a strong indication that the interaction during passage of the wave groups in Experiment 1, which effectively lasts for approximately $T_{\mathrm{int}}\sim 7$ wave periods, is far from enough to reach the quasi-equilibrium state: the ratio $|\Delta U^{\prime}(z)/u_{s}^{\prime}(z)|$ only reaches approximately $0.5$ as figure 3c shows, whereas we shall see in the next section (in turn backed by experiments) that in the final quasi-equilibrium state this fraction is $\gtrsim 1$. The wave field, and consequently the Stokes drift, changes slowly, so $\partial_{t}{\boldsymbol{u}}_{s}$ is negligible. The flow is assumed to be unforced, without, e.g., wind stress, and there is no influence from buoyancy or the Coriolis force. The details of the dissipation are not important to the argument, so we shall not consider them beyond assuming that viscous decay is slow compared with the mean-flow effects of wave–turbulence interactions (which is reasonable in light of the observations by Jooss et al., 2021), so that a quasi-steady state can be reached. The final assumption is that $\bar{{\boldsymbol{u}}}$ and ${\boldsymbol{u}}_{s}$ are both oriented along the $x$-axis and vary slowly except with coordinate $z$.
@@
$\partial_{t}\check{{\boldsymbol{u}}}+(\check{{\boldsymbol{u}}}\cdot\nabla)\check{{\boldsymbol{u}}}={\boldsymbol{u}}_{s}\times\check{\boldsymbol{\omega}}-\nabla(\check{p}+{\textstyle\frac{1}{2}}u_{s}^{2}+{\boldsymbol{u}}_{s}\cdot\check{{\boldsymbol{u}}})+\text{diffusion,}$ (9)
where $\check{\boldsymbol{\omega}}=\nabla\times\check{{\boldsymbol{u}}}=(0,\partial_{z}\bar{u},0)+\nabla\times{\boldsymbol{u}}^{\prime}$ denotes the (Eulerian) wave-averaged vorticity field. Taking the $x$-component and ignoring diffusion gives
@@
Equation (11) has two important consequences. First, that an inhomogeneous field of turbulence will drive a mean current for as long as a vertical gradient of the shear stress exists (and dominates over viscous forces). Second, when a steady state has been reached, i.e., $\partial_{t}\bar{u}=0$, then $\overline{u^{\prime}w^{\prime}}$ is approximately constant with respect to depth (see also Pearson, 2018, for more discussion), or, more precisely, its vertical gradient is balanced by viscous diffusion. Physically, $\overline{u^{\prime}w^{\prime}}\neq 0$ represents a vertical redistribution of streamwise momentum, allowing ${\boldsymbol{u}}={\boldsymbol{u}}(z)$ to transition from its initial value to a presumed final steady-state value, so when the transition period is ended, this shear stress should thus be small compared with TKE. Equation (11) paves the way for further analysis of the ‘spin-up’ of the Eulerian-mean current, because the right-hand side can be related to the underlying physical process using a model based on RDT.
@@ 4.3 Statistics in the quasi-equilibrium state
where we have employed the chain rule and the fact that $\nabla\cdot{\boldsymbol{u}}^{\prime}=0$. The diffusion term contains viscous dissipation and turbulent pressure fluctuations, both of which may reasonably be assumed to be small in an oceanographic setting (Harcourt, 2015; Pearson, 2018) (unfortunately we cannot directly ascertain how accurate this assumption is in the present context; see also Pearson et al., 2019). We should bear in mind that this assumption is questionable for our experiment, where waves are shorter and turbulence levels considerably higher than in typical field settings. This is especially true for cases 2.A.2 and 2.B.2. For instance, angular diffusion of the waves can be highly significant when the waves are not long compared with the turbulent integral scale as found by Smeltzer et al. (2023) which would cause a spatially variable Stokes drift. The approximation is more justified the larger the Stokes drift ($\sim(ak_{0})^{2}c(k_{0})$) is compared with turbulent velocities ($\sim u_{\mathrm{rms}}$). In cases 2.A.2 and 2.B.2 the turbulence is strong and the waves short, and turbulent diffusion and pressure correlation terms may not be negligible compared with the Stokes drift contributions even at surface level. Unfortunately, we are not in a position to quantify the diffusive terms in (12) from our data.
@@
which is to say that an Eulerian-mean current must have been created with the opposite sign compared with Stokes drift.
@@
with ${\boldsymbol{u}}_{s}$ and $\bar{{\boldsymbol{u}}}_{h}$ lying in the horizontal plane. We emphasise that if a constant, depth-uniform current is initially present (as in our experiment), the current $\bar{{\boldsymbol{u}}}$ in (13) and (14) is the change in current due to wave–current interaction, or alternatively the current measured in the reference system following the original Eulerian-mean flow.
Beneath a free surface, turbulence is not isotropic, yet $\overline{u^{\prime}u^{\prime}}$ and $\overline{w^{\prime}w^{\prime}}$ can be expected to be of the same order of magnitude except very near the mean surface level where $\overline{w^{\prime}w^{\prime}}$ tends to zero. It is worth noticing that (13) is not a perfect cancellation between Eulerian flow and Stokes drift as suggested by the experimental (re)analysis of Monismith et al. (2007), but when turbulence is close to isotropic, the remaining Lagrangian-mean current could be difficult to distinguish from zero in an experiment. On the basis of available evidence it seems a reasonable conjecture that the wave–turbulence-generated anti-Stokes flow explains, at least partly, this surprising cancellation.
@@ 4.3.1 Comparison with experiment
Figure 6: Test of equation (13) for cases 2.A to 3.B. The theoretical Stokes drift gradient $\mathrm{d}u_{s}/\mathrm{d}z$ (the derivative of equation 1) is shown as dashed lines of corresponding colour. Separation of turbulence from waves was performed with POD, discussed in Appendix C.
In order to test (13), it is necessary to extract the variances $\overline{u^{\prime}u^{\prime}}$ and $\overline{w^{\prime}w^{\prime}}$ from the wave–turbulence PIV data. Accomplishing this with the best possible accuracy is paramount in order to estimate turbulent second-order moments, because the wave velocities far exceed those of the turbulence near the surface, so even a small percentage of the wave motion erroneously identified as turbulence could lead to considerable overestimates of $\overline{u^{\prime}u^{\prime}}$ and $\overline{u^{\prime}w^{\prime}}$ (less so $\overline{u^{\prime}w^{\prime}}$ due to the phase difference between $\tilde{u}$ and $\tilde{w}$).
@@
The two sides of (13) are compared in figure 6 for cases 2.A to 3.B, the cases whose measurements were taken in the streamwise–vertical plane so that waves and turbulence can be separated. The POD procedure was employed for triple decomposition to evaluate turbulent second moments. The dashed lines show the vertical derivative of the Stokes drift, $\mathrm{d}u_{s}/\mathrm{d}z=2\sqrt{gk_{0}}(ak_{0})^{2}\exp(2k_{0}z)$, while the solid line is the measured value of $(\overline{w^{\prime}w^{\prime}}/\overline{u^{\prime}u^{\prime}})\mathrm{d}\bar{u}/\mathrm{d}z$, which is hypothesised to equal the Stokes drift gradient in the quasi-steady state. For completeness, the Reynolds stresses themselves are plotted in Appendix D.
@@
Given the highly irregular and unsteady nature of our strongly turbulent flow, the agreement is striking in many of the cases. As argued above, the assumptions behind (13) are expected to be reasonable when Stokes drift is sufficiently high compared with turbulent velocities, i.e. $k_{0}z\gtrsim-2$ and sufficiently large $ak_{0}$, and subject to corrections if diffusion levels are high. The relatively poor fit for cases 2.A.2 and 2.B.2 is therefore not very surprising for the latter reason. The very low values of $ak_{0}$ in cases 2.A.1.1 and 2.A.2.1 is a likely reason why (13) is unlikely to be reasonable for these cases – again the Stokes drift term in (25) may not dominate over terms which are neglected. Indeed, the good performance in case 3.A.1 is more striking, but tallies with the fact that TKE is only half that of case 2.B and less than a quarter of that of case 2.A, with correspondingly less diffusion.
Some further words of caution are warranted. As discussed previously, turbulence was measured only at one position near the downstream end of the channel, while the wave–current interaction occurred upstream. We therefore cannot guarantee that the ratio $\overline{u^{\prime}u^{\prime}}/\overline{w^{\prime}w^{\prime}}$ has been the same throughout the time of interaction – indeed the theory of Teixeira & Belcher (2002) indicates this ratio has increased in magnitude near the surface during the spin-up period due to wave–current interaction. We hesitate to offer an explanation for the surprising fact that the experimental curves in panel figure 6 appear to curve back near the surface, corresponding to a similar behaviour of the fraction $\overline{u^{\prime}u^{\prime}}/\overline{w^{\prime}w^{\prime}}$ (see also figure 12(i)), yet we remind the reader of the difficulty of separating waves and turbulence near the surface for these cases; we cannot rule out that the effect is spurious.
@@ 4.4 A model based on RDT
The early onset of the (change in the) Eulerian current is governed approximately by (11), and we will use RDT to study how its right-hand side—that is, the Reynolds stress $\overline{u^{\prime}w^{\prime}}$—depends on depth $z$ and time while changes are still moderate. The theoretical predictions will be compared with the observations made for cases 1.A1.D, shown in figure 3(b). Teixeira & Belcher (2002) showed using RDT that a shear stress that varies in the vertical is generated by the passage of a progressive monochromatic surface wave over isotropic turbulence. We will show next that something similar occurs for a finite wave group, such as considered in the experiments of the present study. Since the effect we consider is inviscid in nature, we will neglect viscosity in the following.
@@
The RDT is a theory first proposed by Batchelor & Proudman (1954) where the straining of turbulence due to the distortions from the surrounding flow is assumed to dominate over that due to turbulence acting on itself, so that the term $({\boldsymbol{u}}^{\prime}\cdot\nabla){\boldsymbol{u}}^{\prime}$ is negligible in the Navier–Stokes equation, yielding a linearised theory. The leading cause of change for the turbulence is presumed to be the mean Lagrangian motion of turbulence-containing parcels of fluid due to the larger-scale surrounding motion. The RDT approach is formally valid whenever the distortions applied to turbulence are sudden. This amounts to assuming that the distortions (for example, mean-flow gradients) are applied over a time shorter than an eddy turn-over time. The spectral formulation of RDT also requires that there is a spatial scale separation between the turbulence and the mean flow. In the present wave-associated mean flow, this corresponds to $\lambda\gg L$ where $\lambda$ is the wavelength of the waves and $L$ is the integral length scale of the turbulence (cf. Teixeira & Belcher (2002)). Both of these criteria are reasonably well satisfied in the experiments 1.A1.D where the relevant turbulent lengthscale, $L_{x}^{x}$ is at most half a wavelength. Visual inspection of the velocity field clearly shows that the typical coherent eddies are much smaller than $\lambda_{0}$. The values of $L_{x}^{x}$ and details of its calculation with the method of Mora & Obligado (2020) were reported in Smeltzer et al. (2023), and are quoted in Table 2. In practice, RDT is known to provide useful results even when these conditions are not strictly fulfilled (e.g. Hunt & Carruthers, 1990; Mann, 1994; Cambon & Scott, 1999).
@@
In the present case, the effect that is essential in order to explain the generation of a Eulerian-mean current is not related to the individual wave oscillations but rather to the systematic tilting and stretching of the vorticity of the turbulence by the Stokes drift of the wave. Therefore, we adopt a linearised version of the Craik–Leibovich equation, (9) whose key term is the ‘vortex force’ ${\boldsymbol{u}}_{s}\times\check{\boldsymbol{\omega}}$. Applying a simplified phase-averaged formalism is a good approximation because individual wave cycles have little effect on the net stretching and tilting of vortices, and it is the integrated effect of the whole wave group which matters; this is illustrated, e.g., in figure 3 of Teixeira & Belcher (2010) (regular waves are considered there, but the arguments hold equally for the relatively long wave groups we consider here). Only a small uncertainty is thereby introduced which is far smaller than the quantitative effects of the simplifications underlying RDT itself. The corresponding vorticity equation, which will be used in RDT, is
@@
In (15), ${\boldsymbol{\omega}}^{\prime}=\nabla\times{\boldsymbol{u}}^{\prime}$ is the turbulent vorticity. As before, ${\boldsymbol{u}}_{s}$ is orientated along the $x$ axis and is independent of $x,y$ and $t$. The mean background current is assumed to be initially depth uniform; hence, it has only trivial effect on the flow system, and we set it to zero by a change of coordinate system. It should be noted that these equations represent the effect of the Stokes drift of an irrotational surface wave; the rotational correction to the wave motion due to the small change in Eulerian current (see Ellingsen, 2016) is neglected.
Let us first consider the turbulence away from the air-water interface, which to a first approximation can be considered homogeneous and isotropic, being denoted by the superscript $(H)$. In this section, it is useful to let the subscript $i=1,2,3$ denote the component of a vector or tensor along directions $x,y$ and $z$, respectively. The turbulent, homogeneous (or bulk) vorticity ${\boldsymbol{\omega}}^{\prime(H)}$ is expressed as a three-dimensional Fourier integral, as
@@
where the hat denotes a Fourier transformed turbulent perturbation quantity and ${\boldsymbol{\kappa}}=(\kappa_{1},\kappa_{2},\kappa_{3})$ is the wavenumber vector.
@@
$\frac{\mathrm{d}\hat{\omega}^{(H)}_{i}}{\mathrm{d}t}$ $=\frac{\partial u_{si}}{\partial x_{j}}\hat{\omega}^{(H)}_{j}$, (17)
@@
Since $u_{si}=u_{s}\delta_{1i}$, (17)-(18) reduce to
@@
where the subscript ‘0’ applied to a variable denotes its value at the initial time, before the turbulence is distorted. It is convenient to define
@@
where $\varepsilon_{ijl}$ is the Levi-Civita permutation symbol and $\kappa=|{\boldsymbol{\kappa}}|$.
We wish to relate the turbulent velocity after distortion by the Stokes drift to the velocity of the initial undistorted turbulence. The Fourier transform of the velocity of the homogeneous turbulence after distortion by the Stokes drift can be related to the corresponding Fourier transform of the vorticity using (23). In turn, the vorticity of the homogeneous turbulence after distortion can be related to the initial turbulent vorticity using (20a)-(20b). Finally, the initial turbulent vorticity can be related to the initial turbulent velocity through